Plz some1 help me to solve this problem:
An 8 litre cylinder contains a mixture of oxygen
and nitrogen, the volume of oxygen being 16% of
the total volume. A few litres of the mixture is
released and an equal amount of nitrogen is added.
Then the same amount of the mixture as before is
released and replaced by nitrogen for the second
time. As a result, the oxygen content becomes 9% of
the total volume. How many litres of mixture is
released each time?

Found 2 solutions by lynnlo, ankor@dixie-net.com:
Answer by lynnlo(4176)
Answer by ankor@dixie-net.com(22740)

An 8 litre cylinder contains a mixture of oxygen
and nitrogen, the volume of oxygen being 16% of
the total volume. A few litres of the mixture is
released and an equal amount of nitrogen is added.
Then the same amount of the mixture as before is
released and replaced by nitrogen for the second
time. As a result, the oxygen content becomes 9% of
the total volume. How many litres of mixture is
released each time?
:
let x = amt of mixture released each time, also the amt of nitrogen added
:
We know that if the mixture is 16% oxygen, it is 84% nitrogen
then 9% oxygen means we have 91% nitrogen
:
Write the decimal equation in the terms of nitrogen
.84(8-2x) + 2x = .91(8)
6.72 - 1.68x + 2x = 7.28
-1.28x + 2x = 7.28 - 6.72
.32x = .56
x = .56/.32
x = 1.75 liters of the mixture is released each time.